# Technical Report on the Checkfor.ai AI-Generated Text Classifier

### v2

@@ Abstract
We present the Checkfor.ai text classifier, a transformer-based neural network trained to distinguish text written by large language models from text written by humans. Checkfor.ai outperforms zero-shot methods such as DetectGPT as well as leading commercial AI detection tools with over 9 times lower error rates on a comprehensive benchmark comprised of ten text domains (student writing, creative writing, scientific writing, books, encyclopedias, news, email, scientific papers, short-form Q&A) and 8 open- and closed-source large language models. We propose a training algorithm, hard negative mining with synthetic mirrors, that enables our classifier to achieve orders of magnitude lower false positive rates on high-data domains such as reviews. Finally, we show that Checkfor.ai is not biased against nonnative English speakers and generalizes to domains and models unseen during training.
@@ 2 Algorithm
Figure 1: Training process for the Checkfor.ai AI-generated text classifier. An initial classifier predicts on a large training pool of human examples, identifying false positives which are then added to the training set and mirrored by LLMs.
@@ 3.1 Overview
Figure 2: Overall results. (Left): Accuracy by detection method. (Right): False positive and false negative rates by detection method. Checkfor.ai has significantly higher accuracy than the next best methods, demonstrating state-of-the-art performance.
@@ 3.2 Overall Performance
Checkfor.ai’s text classifier is the only model that achieves production-ready levels of accuracy, false positive rate, and false negative rate. Our model is the most accurate at 99%, compared to commercial competitors which do not even clear 95%. Our false positive rate is better than the second best model, GPTZero, by a factor of 3, which achieving 7 times better negative error rate.
@@
Figure 3: Accuracy by Domain. Checkfor.ai outperforms GPTZero and Originality on all 10 domains tested, demonstrating robustness to a wide variety of writing styles and formats.
@@ 3.3 Performance by Domain
We compare the performance of Checkfor.ai against GPTZero and Originality on 10 domains of text: blog posts, books, creative writing, email, news, Q&A, reviews, scientific writing, student writing, and Wikipedia. Creative writing is taken mostly from the Reddit WritingPrompts subreddit (Fan et al., 2018) as well as the Ghostbusters subset (Verma et al., 2023). Books are taken from Project Gutenberg. Email is taken from the Enron email dataset (Enron Corp and Cohen, 2015). QA is taken from a mix of sources, including Reddit ELI5, AskHistorians, AskScience (Fan et al., 2019), as well as the open source MedQUAD (Asma and Demner-Fushman, 2019) and FIQA datasets (fiq, 2018). Student writing is taken from a mix of sources, including the PERSUADE corpus (University et al., 2023), the Liang benchmark (Liang et al., 2023), and college admissions essays collected from various sources around the web. Reviews are taken from Amazon, Google Maps, and Yelp directly. In all cases, human data is verified to have been written prior to 2021 to avoid contamination from AI writing.
Figure 4: False Positive and False Negative Rates by Domain. Other models show bias towards over- or under-predicting AI labels. Checkfor.ai is the only model that achieves both low FPR and FNR.
@@ 3.4 Performance by LLM
Figure 5: Recall at 1% FPR by LLM that generated the AI text. Checkfor.ai’s performance remains strong on the most capable model, GPT-4, while the other models experience a severe degradation in performance.
We observe that Checkfor.ai is the only model that can achieve greater than 97% recall on all the LLMs tested. Interestingly, GPTZero performs best on the open-source models. This is likely due to the fact that perplexity and burstiness features can be used from the same model that generated the output– as exact computation of perplexity can only be done in the white-box setting with an open source model.
GPTZero and Originality completely fail when asked to detect AI content produced by GPT-4, both the latest full version of GPT-4 as well as GPT-4-Turbo. According to (Chakraborty et al., 2023), the GPT-4 family of models is the most difficult family to classify based on probability-based features, showing the limitations of the probability-based approaches to detect increasingly capable LLM output. This is the reason why Checkfor.ai uses standard deep learning techniques – so that the model can learn the underlying patterns of speech and voice created by advanced LLMs such as GPT-4, rather than relying on probability estimates from less capable models. Additionally, the recent release of GPT-4 may mean that it was not incorporated into the training sets of the other models.
@@ 3.5 Performance on Nonnative English (ESL)
Figure 6: False positive rates on the TOEFL benchmark. Checkfor.ai is the only model to achieve zero false positive rate on this English as a Foreign Language benchmark.
@@
We hold out the 91 TOEFL essays from our training set and evaluate Checkfor.ai on the benchmark. Due to the composition of our training set, we report a false positive rate of 0% on the TOEFL benchmark. We report results in Figure 6.
@@ 3.6 Performance on Out-of-Distribution Examples
Our final model achieves a false positive rate of 0.8% on the Enron email dataset after hard negative mining, compared to our competitors (who may or may not have email in their training sets) which demonstrate a FPR of at least 2%. After generating AI examples based on the Enron emails, we find that our false negative rate is around 2%. We show an overall accuracy of 98% compared to GPTZero and Originality which perform at 89% and 91% respectively.
@@ 3.7 Performance on LLMs unseen during training
We evaluate our model on two new models released by OpenAI, GPT-4-0125-Preview as well as GPT-3.5-Turbo-0125. We note that the classifier has not seen a single example in the training set that comes from either of these two models. We create a benchmark dataset of 1830 and 1781 examples respectively and evaluate our model on the same metric as above, recall at 1% precision using the same threshold as in Section 2.4. For this experiment, we cannot fairly compare our model to our competitors in this setting, as we do not know whether or not those models have seen these during training.
We find that our model maintains strong performance on these two new models in Table 2, showing robustness to new OpenAI releases.
LLM Recall at 1% FPR
GPT-3.5-Turbo-0125 99.66%
GPT-4-0125-Preview 99.18%
Table 2: Checkfor.ai performance on OpenAI models unseen during training. Our model shows robustness to model updates without needing to regenerate new training data.
We are aware that we won’t be able to train on the outputs of every model or even model family. However, our classifier must perform well on new and out-of-distribution open source models if we expect to serve as an effective safeguard for the internet. We evaluate the Checkfor.ai classifier on evaluation sets of 1300 examples from each of several models available in the Together AI API. The results in Table 3 show that our model effectively generalizes to open source models, with greater than 99.6% recall on all models tested.
@@
Table 3: Checkfor.ai performance on open source models unseen during training. Our model shows robustness to new models and model families trained on open source data.
@@ 4.4 Hard Negative Mining with Synthetic Mirrors
Creative Writing 1.51% 0.19%
Reviews 1.81% 0.09%
Books 0.85% 0.001%
Scientific Papers 1.54% 0.21%
Wikipedia 5.34% 0.23%
News 0.55% 0.04%
Q&A 2.52% 0.13%
ESL 1.44% 0.05%
Student Writing (No Hard Negative Mining) 0.75% 0.75%
Email (Not in training set) 6.60% 0.80%
Total domain-weighted FPR 2.29% 0.25%
@@ 6 Conclusion
In this technical report we have detailed the performance of the Checkfor.ai text classifier against other commercial AI detection systems. We have shown that a simple algorithm for hard negative mining combined with a synthetic mirroring strategy is necessary to scale the performance of a deep-learning based classifier to industry-ready levels of accuracy. We show that the algorithm’s novel ability to scale to web-sized datasets enables generalization to out-of-domain text, LLM-generated text where the LLM has not been observed during training, and ability to detect AI-generated text from even the current most capable model, GPT-4.
@@ 7 Ethics and Responsible Use
Our classifier is available for public commercial use at https://checkfor.ai. We strongly discourage the use of our classifier as a sole arbiter of academic integrity and plagiarism checking. All AI detection tools have a nonzero false positive rate, and should be used in conjunction with other evidence to prove or disprove plagiarism.

### v3

# Technical Report on the Pangram AI-Generated Text Classifier

@@ Abstract
We present Pangram Text, a transformer-based neural network trained to distinguish text written by large language models from text written by humans. Pangram Text outperforms zero-shot methods such as DetectGPT as well as leading commercial AI detection tools with over 38 times lower error rates on a comprehensive benchmark comprised of 10 text domains (student writing, creative writing, scientific writing, books, encyclopedias, news, email, scientific papers, short-form Q&A) and 8 open- and closed-source large language models. We propose a training algorithm, hard negative mining with synthetic mirrors, that enables our classifier to achieve orders of magnitude lower false positive rates on high-data domains such as reviews. Finally, we show that Pangram Text is not biased against nonnative English speakers and generalizes to domains and models unseen during training.
@@ 2 Algorithm
Figure 1: Training process for the Pangram AI-generated text classifier. An initial classifier predicts on a large training pool of human examples, identifying false positives which are then added to the training set and mirrored by LLMs.
@@ 3.1 Overview
Figure 2: Overall results. (Left): Accuracy by detection method. (Right): False positive and false negative rates by detection method. Pangram has significantly higher accuracy than the next best methods, demonstrating state-of-the-art performance.
@@ 3.2 Overall Performance
Pangram’s text classifier is the only model that achieves production-ready levels of accuracy, false positive rate, and false negative rate. Our model is the most accurate at 99%, compared to commercial competitors which do not even clear 95%. Our false positive rate is better than the second best model, GPTZero, by a factor of 3, which achieving 7 times better negative error rate.
@@
Figure 3: Accuracy by Domain. Pangram outperforms GPTZero and Originality on all 10 domains tested, demonstrating robustness to a wide variety of writing styles and formats.
@@ 3.3 Performance by Domain
We compare the performance of Pangram against GPTZero and Originality on 10 domains of text: blog posts, books, creative writing, email, news, Q&A, reviews, scientific writing, student writing, and Wikipedia. Creative writing is taken mostly from the Reddit WritingPrompts subreddit (Fan et al., 2018) as well as the Ghostbusters subset (Verma et al., 2023). Books are taken from Project Gutenberg. Email is taken from the Enron email dataset (Enron Corp and Cohen, 2015). QA is taken from a mix of sources, including Reddit ELI5, AskHistorians, AskScience (Fan et al., 2019), as well as the open source MedQUAD (Asma and Demner-Fushman, 2019) and FIQA datasets (fiq, 2018). Student writing is taken from a mix of sources, including the PERSUADE corpus (University et al., 2023), the Liang benchmark (Liang et al., 2023), and college admissions essays collected from various sources around the web. Reviews are taken from Amazon, Google Maps, and Yelp directly. In all cases, human data is verified to have been written prior to 2021 to avoid contamination from AI writing.
Figure 4: False Positive and False Negative Rates by Domain. Other models show bias towards over- or under-predicting AI labels. Pangram is the only model that achieves both low FPR and FNR.
@@ 3.4 Performance by LLM
Figure 5: Recall at 1% FPR by LLM that generated the AI text. Pangram’s performance remains strong on the most capable model, GPT-4, while the other models experience a severe degradation in performance.
We observe that Pangram is the only model that can achieve greater than 97% recall on all the LLMs tested. Interestingly, GPTZero performs best on the open-source models. This is likely due to the fact that perplexity and burstiness features can be used from the same model that generated the output– as exact computation of perplexity can only be done in the white-box setting with an open source model.
GPTZero and Originality completely fail when asked to detect AI content produced by GPT-4, both the latest full version of GPT-4 as well as GPT-4-Turbo. According to (Chakraborty et al., 2023), the GPT-4 family of models is the most difficult family to classify based on probability-based features, showing the limitations of the probability-based approaches to detect increasingly capable LLM output. This is the reason why Pangram uses standard deep learning techniques – so that the model can learn the underlying patterns of speech and voice created by advanced LLMs such as GPT-4, rather than relying on probability estimates from less capable models. Additionally, the recent release of GPT-4 may mean that it was not incorporated into the training sets of the other models.
@@ 3.5 Performance on Nonnative English (ESL)
Figure 6: False positive rates on the TOEFL benchmark. Pangram is the only model to achieve zero false positive rate on this English as a Foreign Language benchmark.
@@
We hold out the 91 TOEFL essays from our training set and evaluate Pangram on the benchmark. Due to the composition of our training set, we report a false positive rate of 0% on the TOEFL benchmark. We report results in Figure 6.
@@
We further evaluate Pangram Text on two additional ESL datasets.
We evaluate Pangram on 3,907 essays from the ELLIPSE dataset Franklin et al. (2022), which contains ESL essays from 8th-12th graders, and find zero false positives.
We also evaluate Pangram on the recently released ICNALE dataset Ishikawa (2023), which contains 5,600 essays from undergraduates at Asian universities learning English. We report a false positive rate of 0.09% on this dataset.
Based on these findings, we conclude that Pangram is not biased against text written by non-native English speakers.
@@ 3.6 Performance on Out-of-Distribution Examples
Our model with email held out achieves a false positive rate of 0.8% on the Enron email dataset after hard negative mining, compared to our competitors (who may or may not have email in their training sets) which demonstrate a FPR of at least 2%. After generating AI examples based on the Enron emails, we find that our false negative rate is around 2%. We show an overall accuracy of 98% compared to GPTZero and Originality which perform at 89% and 91% respectively.
After retraining on all Enron email, except for a test set of 1000 held-out Enron email documents, we are able to achieve zero errors on the test set.
@@ 3.7 Performance on LLMs unseen during training
We are aware that we won’t be able to train on the outputs of every model or even model family. However, our classifier must perform well on new and out-of-distribution open source models if we expect to serve as an effective safeguard for the internet. We evaluate the Pangram classifier on evaluation sets of 1300 examples from each of several models available in the Together AI API. The results in Table 2 show that our model effectively generalizes to open source models, with greater than 99.6% recall on all models tested.
@@
Table 2: Pangram performance on open source models unseen during training. Our model shows robustness to new models and model families trained on open source data.
3.8 July 2024 Update: Performance on Recently Released LLMs
We evaluated Pangram Text on the major recently released commercial large language models, as of July 2024, using a benchmark of 25,000 documents generated in the same way as the original benchmark. We observe that we are able to continue achieve excellent performance on these newly released models.
LLM Recall at 1% FPR
GPT 4o 100.0%
Claude 3 (various sizes) 99.76%
LLaMA 3 (various sizes) 99.97%
Table 3: Pangram performance on recently released models from July 2024.
3.9 July 2024 Update: Performance on non-English Languages
In July 2024 we updated Pangram text to distinguish multilingual human and AI-generated text. To benchmark our performance, we use the Amazon Multilingual Reviews dataset (Keung et al., 2020), the XLSum dataset (Hasan et al., 2021), and the multilingual Wikipedia datasets, and generate synthetic mirrors for each.
Language Amazon Reviews Accuracy Wikipedia Accuracy XLSum (BBC News) Accuracy
Spanish 99.59% 99.75% 99.75%
French 98.84% 99.33% 98.50%
Italian N/A 99.82% N/A
German 99.44% 99.95% N/A
Portuguese N/A 99.83% 99.70%
Russian N/A 98.34% 99.35%
Chinese 99.70% 99.54% 98.10%
Table 4: Pangram multilingual performance.
@@ 4.4 Hard Negative Mining with Synthetic Mirrors
Creative Writing 1.51% 0.02%
Reviews 1.81% 0.02%
Books 0.85% 0.01%
Scientific Papers 1.54% 0.04%
Wikipedia 5.34% 0.05%
News 0.55% 0.001%
Q&A 2.52% 0.009%
ESL 1.44% 0.01%
Student Writing (No Hard Negative Mining) 0.75% 0.04%
Email 6.60% 0.00%
Total domain-weighted FPR 2.29% 0.02%
@@ 6 Conclusion
In this technical report we have detailed the performance of the Pangram text classifier against other commercial AI detection systems. We have shown that a simple algorithm for hard negative mining combined with a synthetic mirroring strategy is necessary to scale the performance of a deep-learning based classifier to industry-ready levels of accuracy. We show that the algorithm’s novel ability to scale to web-sized datasets enables generalization to out-of-domain text, LLM-generated text where the LLM has not been observed during training, and ability to detect AI-generated text from even the current most capable model, GPT-4.
@@ 7 Ethics and Responsible Use
Our classifier is available for public commercial use at https://pangram.com. We strongly discourage the use of our classifier as a sole arbiter of academic integrity and plagiarism checking. All AI detection tools have a nonzero false positive rate, and should be used in conjunction with other evidence to prove or disprove plagiarism.